Please help me on these type of questions:
Tell whether the graphs of each pair of equations are parallel or perpendicular.
1.) 6x+5y=5 2.) 2x+4y=8 3.) 2x+y=4
6x+5y=10 y=2x-8 y=-2x+1
I don't get this. As in I don't know how to tell if they are parallel or perpen. and to graph them. Please help me!

Answer by checkley71(8403)

Y=mX+b IS THE STANDARD LINE EQUATION.
THUS YOU HAVE TO PUT EACH EQUATION IN THIS FORM AND COMPARE THE m VALUES.
IF THEY ARE EQUAL THEN THEY ARE PARALLEL.
IF THEY ARE THE NEGATIVE RECIPRICAL THEN THEY ARE PERPENDICULAR. 2 & -1/2 ARE NEGATIVE RECI[PRICALS OF EACH OTHER.
6X+5Y=5 OR 5Y=-6X+5 OR Y=-6X/5+5/5 OR Y=-6X/5+1 SLOPE=(-6/5)
6X+5Y=10 OR 5Y=-6X+10 OR Y=-6X/5+10/5 OR Y=-6X/5+2 SLOPE=(-6/5)
THEREFORE THESE 2 LINE SARE PARALLEL.
------------------------------------------------------------------
2X+4Y=8 OR 4Y=-2X+8 OR Y=-2X/4+8/2 OR Y=-X/2+4 SLOPE=(-1/2)
Y=2X-8 SLOPE=(2) THUS THESE 2 LINES ARE PERPENDICULAR.
------------------------------------------------------------------
I'LL LEAVE THE THIRD PROBLEM FOR YOU TO PRACTICE WITH.
GOOD LUCK : )